GL901 - Based on Problem 9-1A LO C2, P1

The January 1, Year 1 trial balance for the Carter Company is found on the trial balance tab. The beginning balances are assumed.

Gonzalez Co. entered into the following transactions involving short-term liabilities. (Use 360 days a year.)

Year 1
Apr. 20 Purchased $49,750 of merchandise on credit from Nguyen, terms n/30.
May 19 Replaced the April 20 account payable to Nguyen with a 90-day, 8%, $38,000 note payable along with paying $11,750 in cash.
July 8 Borrowed $102,000 cash from NMR Bank by signing a 120-day, 10%, $102,000 note payable.
Aug. 17 Paid the amount due on the note to Nguyen at the maturity date.
Nov. 5 Paid the amount due on the note to NMR Bank at the maturity date.
Nov. 28 Borrowed $60,000 cash from Chicago Bank by signing a 60-day, 6%, $60,000 note payable.
Dec. 31 Recorded an adjusting entry for accrued interest on the note to Chicago Bank.
Year 2
Jan. 27

Paid the amount due on the note to Chicago Bank at the maturity date.

www.uy....wwwwwwwwwwwIVI..vwwww.................... The January 1, Year 1 trial balance for the Carter Company is found on thJoll. 21 Poru Lie IUUIL Uue UI Lie lule LU LILILagu Dark ALLiela LUI ILYUule. Answer is not complete. Requirement General JouAnswer is not complete. Requirement General Journal General Ledger Trial Balance Schedule of Payables Calculation of InterestDel. 51 Kelurueu all du jusling ellery TUI dllrueu Interest UHLE HULE LU CHIldgu balik. Year Jan. 27 Paid the amount due on tRequirement General Journal General Ledger Trial Balance Schedule of Payables Calculation of Interest Year 2 payment Enter th
Answer #1
Upto the Stage 4-Trial Balance all are correct.
Stage 5: Schedule of Payables:
Gonzalez Co.
Notes Payable:
Notes payable - chicago 60000
Total Notes payable 60000
Interest Payable:
Interest Payable - Chicago 330
Total Interest Payable 330
Stage 6: Calculation of Interest:
Principal = 60000
Period = 60
Rate of interest = 6%
Number of days in a year = 360
Total interest for 60 days = 60000*6%*60/360 =600
Interest payable on Dec 31,2018 = 60000*6%*33/360 =330
Interest payable in Year 2 = 600-330=270
Stage 7: Year 2 payment entry:
Date Acc. Titles Debit $ Credit $
1/27/Year 2 Notes payable-Chicago 60000
Interest payable 330
Interest expense 270
Cash 60600
(being Chicago note paid off along with interest)
Answer #2

Year 2 Payment: